OpenStudy (anonymous):

how do you graph y=x^2-4x+7?

OpenStudy (anonymous):

First you have to find the line of reflection, since quad eq. symmetric. 

OpenStudy (anonymous):

In that equation, y = x^2 the line of reflection is x = 0

OpenStudy (anonymous):

The figure I drew, I mean

OpenStudy (anonymous):

like how do i know where to put the points because i need to label the vertex and axis of symmetry

OpenStudy (anonymous):

The axis of symmetry IS the line of reflection. Since x = "Whatever you got for the axis of symmetry", then you can subsitute x for that in the quadratic equation, giving you the value for the y, giving you (x,y), giving you the vertex!

OpenStudy (jdoe0001):

hmm, you know the components a b and c for that, right?

OpenStudy (jdoe0001):

that is for \(\bf ax^2+bx+c\)

OpenStudy (anonymous):

a=x^2 b=4x and c=7

OpenStudy (jdoe0001):

well, just the coefficients only

OpenStudy (anonymous):

i have no idea.... im sorry! im terrible at math.

OpenStudy (anonymous):

what exactly would the coefficiets be?

OpenStudy (anonymous):

a = 1 b = 4 c = 7

OpenStudy (jdoe0001):

\(\large \begin{array}{lllll}\\ y=&1x^2&-4x&+7\\ &a&b&c \end{array}\)

OpenStudy (jdoe0001):

just the coefficients though thus a = 1, b = -4 and c = +7

OpenStudy (anonymous):

ok and then i plug that into the quadratic equation?

OpenStudy (anonymous):

Quadatic equations only find the value of x when y = 0. It won't really do any help when graphing multiple points.

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

so to get the vertex for a parabola like so notice that that the leading coefficient, that is, the "a" component is positive, that means is going upwards anyhow to get the vertex of it, the coordinates of it will be at \(\bf \left(-\cfrac{b}{2a}\quad , \quad c-\cfrac{b^2}{4a}\right)\)

OpenStudy (anonymous):

No, I think it would be (-b/2a, f(-b/2a))

OpenStudy (jdoe0001):

same thing

OpenStudy (anonymous):

and that finds the vertex every time?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Yes, for equations like "ax^2 +bx+c=f(x) or y

OpenStudy (anonymous):

thank you so much! you don't understand how much you've helped! i learn more here than in my actual class. thanks again! :)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

That is the power of Open Study.

OpenStudy (anonymous):

i'll defiantly be coming on here more often haha

OpenStudy (anonymous):

We hope to answer more of your questions!

OpenStudy (anonymous):

pretty sure i'll have more after i finish this one hahaha